Writing and solving two step equations: Translate the statement into an equation. Then solve the equation. Page 136 problem 10
Seven subtracted from the quotient of a number and 2 is -6.

Answer by checkley77(12844)

2x-7=-6
2x=-6+7
2x=1
x=1/2 answer.